# Target Speaker ASR with Whisper

### v1

@@ III-D Scaling System With More Data And Parameters
Given that even a simple approach using a single bias performs well, Table V demonstrates the improvement in system performance when additional training data is used. The results highlight that incorporating synthetic data provides additional gains beyond those achieved with real meeting data alone, raising the question of whether pretraining on synthetic data might offer further improvements.
Table VI provides a performance analysis across different model sizes, highlighting the improvements gained from employing an additional CTC head. The analysis also explores the impact of CTC preheating and an Amplification phase, showing how these techniques can optimize performance across varying model scales.
@@ IV Conclusions and Limitations
In this study, we have demonstrated the efficacy of our approach and analyzed its setup across multi-domain datasets. However, further validation with additional datasets, especially in different conditions or languages, is necessary to confirm its generalizability and scalability.
Our analysis has highlighted both strengths and limitations of the model. Nevertheless, a more in-depth examination of errors and performance under specific conditions is essential to understand and fully optimize the system. Future work will analyze these errors and explore strategies to improve the model’s robustness against diarization errors.
The investigation into the usage of synthetic data suggests that training a TS-ASR model from scratch using only synthetic data with learnable embeddings, followed by fine-tuning on target data, is also a promising avenue. It remains to be seen whether this approach will achieve the same level of performance as models like Whisper, which benefit from extensive pre-training on diverse datasets.
Moreover, our method can be applied to other pre-trained ASR models. Integrating diarization information into these models using our approach could provide valuable insights into its versatility and effectiveness. Comparing the performance of TS-ASR across various ASR architectures will be an important step in evaluating its adaptability and benefits.

### v2

@@ III-D Scaling System With More Data And Parameters
Table V explores the impact of incorporating additional training data on system performance. The results show that adding the AMI dataset improves the performance on the NOTSOFAR-1, highlighting the benefits of more real-world data. Furthermore, incorporating Libri2Mix data provides additional gains, raising the question of whether pretraining on synthetic data could offer even greater performance gains.
Finally, Table VI provides a performance analysis across different model sizes, highlighting the improvements from employing an additional CTC head. Notably, incorporating a randomly initialized CTC head does not improve performance and leads to degradation. However, preheating the CTC head provides some improvements, especially for the small and medium models, even without joint decoding. Additionally, an FDDT preheating phase further enhances performance, with the best results observed in the large-v3 model.
@@ IV Conclusions and Limitations
In this study, we introduced FDDT into the Whisper model, making it a target-speaker ASR (TS-ASR) system conditioned on diarization output. We analyzed factors such as model size, FDDT placement, dataset size, and parameter initialization, evaluating the system on both real and synthetic datasets with consistently strong results. We also examined the impact of incorporating a CTC head, noting some improvements when it was preheated.
While the FDDT-based TS-ASR method performed effectively across diverse datasets, further validation on varied datasets, conditions, and languages is needed. We identified strengths and limitations, particularly regarding robustness against diarization errors, which need further attention for better real-world performance. Finally, the approach is extendable to other pre-trained ASR models, and a comparative analysis across ASR architectures could offer deeper insights into its adaptability and benefits.
Acknowledgement
The work was supported by Ministry of Education, Youth and Sports of the Czech Republic (MoE) through the OP JAK project ”Linguistics, Artificial Intelligence and Language and Speech Technologies: from Research to Applications” (ID:CZ.02.01.01/00/23_020/0008518) and Brno Ph.D. Talent Scholarship Programme. Computing on IT4I supercomputer was supported by MoE through the e-INFRA CZ (ID:90254).